# Fundamental Oscillations of Massive Boson Stars and Distinguishability

### v2

@@ Abstract
Massive Boson Stars are self-gravitating configurations of self-interacting scalar fields. The equation of state of massive boson stars and their masses, radii, modeled by a self-interacting scalar field with potential of the form $V(\phi)=\frac{1}{2}m^{2}|\phi|^{2}+\frac{1}{4}\lambda|\phi|^{4}$ are known to follow scaling relations. The non-radial fundamental oscillations of such massive BSs have been studied only for a few select model parameters so far. In this work, we demonstrate for the first time that the $f$-mode characteristics also follow a scaling in the strong interaction limit ($\lambda\gg m^{2}/M_{Pl}^{2}$). This opens up the outstanding prospect of studying the $f$-modes of massive BSs throughout the scalar DM parameter space. We study the implications of this finding by carrying out a detailed study of massive BS $f$-modes in a separate work. Here, having introduced this scaling, we use it to compare boson star oscillations with the neutron star and black hole quasinormal modes, thus providing a smoking gun for the distinguishability of BSs using gravitational waves.
@@
Boson Stars — Exotic Compact Objects — Oscillations — Gravitational Waves
I Introduction
@@
Dark matter (DM) and its nature are among the most intriguing open questions of physics. DM forms $\sim 25\%$ of the energy budget of our universe. However, its particle nature has remained elusive. DM candidate masses span 90 orders of magnitude from ultralight fuzzy DM to massive compact halo objects (MaCHOs), and primordial black holes (PBHs) [8]. Historically, the most popular candidate for DM has been Weakly Interacting Massive Particles (WIMPs). However, years of experimental effort in the search for WIMPs have been fruitless so far.
On the other hand, the discovery of Higgs Boson has been one of the most important discoveries, theoretically motivated by shortcomings of the standard model (SM) of particle physics. This discovery also established the existence of scalar (spin-0) bosons. Another problem in particle physics, the well-known strong CP violation, has resulted in the prediction of a new scalar particle, axion, which makes a natural candidate for DM (See [9, 15] for a recent review on Axion DM). As a consequence, DM particles such as light bosonic particles like axions, axion-like particles (ALPs), fuzzy DM, etc., have attracted the spotlight lately.
@@
Boson stars (BSs) are stellar configurations of bosonic particles (see [51, 30] for updated review on BSs). These can form from primordial scalar fluctuations, gravothermal collapse in an early matter-dominated era [39], or via a process called gravitational cooling [41]. They can also be formed via trapping of DM in objects like stars and planets [19, 23, 22, 24]. BS is one of the possible exotic compact objects (ECOs) that can mimic other compact objects like black holes (BHs) and neutron stars (NSs). BSs, if they exist, would also contribute to the total DM budget (see [13] for a review on ECOs) of the universe. Their existence would be a key to the DM problem. Thus, observation of BSs can revolutionize our understanding of the universe.
@@
Conventional electromagnetic telescopes can only probe BSs if they are accreting matter [33]. This method is inaccessible in the case of non-accreting BSs. However, the successful detection of gravitational waves [1] has opened up a new window to probe the dark sectors of the universe. Further, the planning of the next-generation detectors covering a wide frequency range makes the study of all possible GW sources of BSs very crucial and timely.
There have been considerable efforts to find observational signatures that will be able to distinguish BSs from BHs.  [12, 42, 25] suggested the distinguishability of BSs and BHs as well as NSs based on tidal deformability. Krishnendu et al. [28, 29] showed that the spin-induced higher multipole moments can be used to distinguish BH from other ECOs/BH mimickers. Later on, Pacilio et al. [34, 48] showed that coherently including the spin-induced quadrupole moment in the waveform along with tidal effects can be used to obtain complementary bounds on bosonic self-interactions with next-generation GW detectors. All these studies are restricted to binary systems and the inspiral phase of mergers. GWs from isolated BSs are again restricted to accreting cases. In this work, we propose a new signature based on $f$-mode oscillations to delineate massive BSs in the strong-interaction limit from BHs and NSs.
@@
To do this, we compute the non-radial quadrupolar fundamental frequency that couples the strongest with gravitational waves. The study of quasinormal modes (QNMs) is crucial as it carries information about the nature of the underlying compact object [10, 11]. There have been a handful of studies on BS QNMs [26, 52, 6, 32][50] is the first work studying $f$-modes for massive BSs in the strong-interaction limit. The study was restricted to NS mass BSs, and the parameters were restricted by the bounds on self-interaction cross section $100.1<\sigma/m<$ 10 [cm2/g]. Another work [14] appeared during the completion of this work that explored select BS models in the same regime.
@@
In this letter, we show for the first time that scaling relations are followed for the $f$-mode frequency and damping time in the strong-interaction limit. This scaling result is significant, as it allows commenting on the $f$-mode characteristics of BSs with arbitrary self-interacting scalar DM parameters within the strong interaction limit. Using this, we make a detailed comparison of BS $l=2$ QNMs with those of BHs and NSs to comment on the distinguishability. The analysis and implications of the scaling, along with the detectability, will be reported in a separate publication [43].
Throughout this work, we use the convention $\hbar=c=1$ and $G=1/M_{Pl}^{2}$, where $M_{Pl}=1.22\times 10^{22}\text{MeV}=2.18\times 10^{-8}\text{kg}$ is the Planck mass.
II Model
@@
For pure Bosonic dark stars, we describe DM by scalar field ($\phi$) self-interacting via quartic self-interactions as described in [17, 50]. The Lagrangian is given by
@@
where $\lambda$ is the self-interaction parameter and $m$ is the DM particle mass. The self-interaction is negligible only if $\lambda|\phi|^{4}/m^{2}|\phi|^{2}\ll 1$ [17], where $|\phi|\sim M_{Pl}$. A dimensionless parameter is defined as
$\Lambda\equiv\frac{\lambda M_{Pl}^{2}}{4\pi m^{2}}~{}.$ (3)
For $\Lambda\gg 1$[17] showed that massive stars can be formed with mass resembling that of a degenerate fermionic star. In this case, $M_{max}=0.22\sqrt{\Lambda}M_{Pl}^{2}/m=0.06\sqrt{\lambda}M_{pl}^{3}/m^{2}$ [17]. Moreover, $\Lambda\gg 1$, can be achieved for $\lambda\ll 1$. In this limit, even if $\lambda\ll 1$, the interaction is important, and the macroscopic properties are dictated by the strength ($\sqrt{\lambda}/m^{2}$).
@@ III Scaling
[17] showed that the energy density and pressure scale as $\rho^{\prime}=\rho x^{2},~{}p^{\prime}=px^{2}$ to obtain the scaled version of the equation of state
@@
The dimensionless mass ($M^{\prime}$) and radius ($R^{\prime}$) can then be defined for BSs as $M^{\prime}=M/(xM_{Pl}^{3})~{},~{}R^{\prime}=R/(xM_{Pl})$ and can be used to obtain scaled TOV equations.
@@
$f$ $=f^{\prime}/(xM_{Pl})~{},$ (8)
$\tau$ $=\tau^{\prime}(xM_{Pl})~{}.$ (9)
We explicitly derive this scaling using the $f$-mode perturbation equations for the case of Cowling approximation (where the background metric is held constant) as well as in full general relativity (see supplementary material).
IV $f$-modes
In this section, we discuss the $f$-mode characteristics of massive BSs in the strong-interaction limit. When the scaling as mentioned in Eq. 7 is used, the $f$-mode equations become completely independent of the parameter $x$ containing the DM parameters (see supplementary material).
@@ IV $f$-modes
Figure 1: The scaled $f$-mode frequency ($f^{\prime}=f(xM_{Pl})$) for massive BSs in the strong interaction limit ($\Lambda\gg 1$) as a function of scaled mass $M^{\prime}=M/(xM_{Pl}^{3})$ and compactness $C$.
Figure 2: The scaled $f$-mode damping time ($\tau^{\prime}=\tau/(xM_{Pl})$) for massive BSs in the strong interaction limit ($\Lambda\gg 1$) as a function of scaled mass $M^{\prime}=M/(xM_{Pl}^{3})$ and compactness $C$.
@@
This single solution is sufficient to reproduce all the results and for the entire parameter space in the strong coupling regime ($\Lambda\gg 1$) using the relations $\tau^{\prime}=\tau/(xM_{Pl})$, $M^{\prime}=M/(xM_{Pl}^{3})$ and $C^{\prime}=C$.
@@
We have performed a detailed analysis of these solutions and also derived analytical fits for these quantities. This will be reported in a separate publication [43].
V Distinguishability/ Comparison with neutron stars and black holes
It would be interesting to compare the BS frequencies with those of the QNMs of other astrophysical compact objects. Here, we compare the fundamental frequency for NSs and BHs, which are dominant modes in the context of GW.
@@ V Distinguishability/ Comparison with neutron stars and black holes
$f_{BS}=f^{\prime}(C)M^{\prime}(C)\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)$ (13)
$\boxed{f_{BS}=202.67f^{\prime}(C)M^{\prime}(C)\left(\frac{M_{\odot}}{M_{BS}}\right)\text{kHz}}~{}.$ (14)
This is a very useful relation as the frequency can be computed using just two physical parameters, i.e., $M$ and $C$. We can use the unique solutions obtained in the previous section to evaluate $f^{\prime}(C)$ and $M^{\prime}(C)$ (see Fig. 1). The analytical fits for these solutions will be provided in [43]. For both cases, the fits are accurate within $0.25\%$ for $M^{\prime}(C)$ and $1\%$ for $f^{\prime}(C)$. Thus, the error in Eq. 13 is less than $1.25\%$ if one uses the proposed analytical relations. The blue regions in Fig. 3 are obtained by varying the compactness for a given mass. The case of maximum compactness (dark blue curve in Fig. 3) is given by
$f_{BS,max}$ $=f^{\prime}(C_{max})M^{\prime}(C_{max})\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)$ (15)
$\approx 0.013\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)~{}.$ (16)
Thus we arrive at the equation,
$\boxed{f_{BS,max}=2.67\left(\frac{M_{\odot}}{M_{BS}}\right)\text{kHz}}~{}.$ (17)
@@
Figure 3: $f$-mode frequency as a function of mass for (a) NS, (b) stellar-mass BH, and (c) SMBH mass range for massive BSs. The black line is the corresponding BH QNM frequency [27]
The $f$-mode oscillations of NSs have been studied in great detail. Given the underlying uncertainty in NS EoS, we obtain a band in the $f$-$M$ space. NSs lie in the mass range of 1-2.5 $M_{\odot}$ with compactness in the range 0.1-0.3. We show this mass range in Fig. 3(a). The blue contour shows the region covered by BSs. The upper boundary corresponds to $C=C_{max}=0.16$. Silver curves show the $f-M$ region spanned by NSs due to the uncertainty in the microscopic equation of state (taken from [37]). We see a partial overlap with the NS band, but BS frequencies in the higher mass range are lower than that of NSs. Thus, GW observation of $f$-mode oscillations of a compact object in the NS mass range but outside the expected NS frequency band would be a signature for BS.
Astrophysical BHs are not expected in the mass range of NSs. However, PBHs, if they exist, may have mass in the solar and sub-solar mass range. On the other hand, stellar mass BHs fall in the range 5-100 $M_{\odot}$ and supermassive BHs (SMBH) in galactic centers have masses $10^{6}$-$10^{9}$ $M_{\odot}$. We show these two mass ranges Figs. 3(b) and 3(c), respectively. BHs with masses larger than 100 $M_{\odot}$ are possible via mergers of lighter stellar mass BHs [2]. We show the $n=0$, $l=2$ fundamental $f$-mode QNM frequency [27] by the black curves in Fig. 3. This is given by $\omega M=0.37+0.19a$ [27] in geometrical units ($G=c=1$), $a$ is the angular momentum, which after reverting to natural units for the non-spinning case means
$f_{BH}$ $=\left(\frac{0.37}{2\pi}\right)\times\frac{1}{GM_{BH}}$
$\approx 0.059\left(\frac{M_{Pl}^{2}}{M_{BH}}\right)$ (18)
$=12.05\left(\frac{M_{\odot}}{M_{BH}}\right)\text{kHz}~{}.$ (19)
The BH frequency only depends on the mass of a non-rotating, chargeless BH [27]. The BH frequency is also consistently higher than that of BS by a certain factor. This is expected from the high compactness of BH ($C=0.5$), which is higher by a factor of $3.125$ compared to the most compact BS. This factor is readily calculable using Eqs. 16 and V. We define a ratio ($\mathcal{R_{BS/BH}}$) as the ratio of the fundamental QNM frequency of the most compact BS to that of BH of the same mass. This ratio is a constant, independent of mass, and is given by
@@
The BH frequency is higher throughout by a factor of $4.50.22^{-1}=$ 4.5. If we observe $f$-modes from dark compact objects consistently with a frequency less than $\mathcal{R_{BS/BH}}$ times the expected BH frequency, it would be a smoking gun for massive BSs. Note that a similar analysis for other potentials ($V(\phi)$), as well as for another type of BSs, e.g., Oscillatons and Proca Stars, is necessary to distinguish between different models for stars made of Bosonic DM. However, this still acts as a clear signature for distinguishing BHs from BSs.
In general, the ratio is given by
$\frac{f_{BS}}{f_{BH}}=16.98f^{\prime}(C)M^{\prime}(C)~{}.$ (21)
A similar relation was provided in [21] for the innermost stable circular orbit (ISCO) frequency ($f_{ISCO}$)
@@
Here, $M$ is the component mass in an equal mass binary. The ISCO frequency marks the end of the inspiral phase during the binary merger. For $C=0.5$, we get the ISCO frequency for BH. We compare the $f$-mode frequency with the ISCO as it marks the importance of dynamical tides during pre-merger in the case of a BS binary system. For BHs, the ratio is
$\frac{f_{BH}}{f_{BH}^{ISCO}}=0.37\times 6^{3/2}=5.44~{}.$ (23)
Note that this number is not as important as the one for BSs, as there are no tidal deformations in the case of binary inspiral BH. For BSs, we get
@@
For the case of BSs, the ratio depends on the compactness. The maximum value obtained for the most compact case is 6.79, which is higher than that for BH. The reason is that although the QNM frequency is higher for BH, the ISCO frequency is much lower in the case of BSs, which are much less compact. It is also a much higher ratio, especially when compared with NSs. The typical ISCO frequency is roughly between 500 Hz - 1kHz [7], and the $f$-mode frequencies lie between 1.5-2 kHz for typical NS masses. Hence, the ratios for NSs is roughly around the range 2-4. Therefore, we expect the dynamical $f$-mode effects to be less prominent in the case of BSs.
VI Summary

### v3

# Fundamental oscillations as a tool to distinguish boson stars from neutron stars and black holes

@@ Abstract
Massive boson stars are self-gravitating configurations of self-interacting scalar fields and can be modeled by a massive scalar field with a quartic self-interaction potential. It has been shown that the equation of state and static structure properties, such as mass and radius, follow scaling relations independent of microscopic dark matter properties. In this work, we demonstrate for the first time that non-radial fundamental ($f$-)mode characteristics also follow a scaling in the strong interaction limit, opening up the outstanding prospect of evaluating the mode properties for boson stars for arbitrary masses spanning the scalar dark matter parameter space allowed by current observations. We provide the scaling relations within full general relativity and obtain the mode characteristics corresponding to the maximum boson star mass configuration. We apply these to determine the $f$-mode properties for boson stars solely as a function of their mass and compactness, which allows distinguishing them from those of neutron stars and black hole quasinormal modes in comparable mass range. In particular, we show that the frequencies are always lower than those of corresponding black holes of the same mass by a factor of 4.5. This provides a smoking gun for the distinguishability of boson stars from other compact objects using gravitational wave observations.
keywords:
@@
Boson Stars , Exotic Compact Objects , Oscillations , Gravitational Waves
journal: Physics of the Dark Universe
[…]
[label3] organization=Institute for Theoretical Physics, Goethe University, addressline=Max von Laue Str. 1, city=Frankfurt am Main, postcode=60438, state=, country=Germany
1 Introduction
@@
Dark matter (DM) and its nature are among the most intriguing open questions of physics. DM forms $\sim 25\%$ of the energy budget of our universe. However, its particle nature has remained elusive. DM candidate masses span 90 orders of magnitude from ultralight fuzzy DM to massive compact halo objects (MaCHOs) and primordial black holes (PBHs) [8]. Historically, the most popular candidate for DM has been Weakly Interacting Massive Particles (WIMPs). However, years of fruitless experimental search for WIMPs, discovery of Higgs Boson, prediction of axions to solve the strong CP violation, failure of cold DM paradigm on smaller scales, etc. has renewed interest in scalar bosonic particles like axions, axion-like particles (ALPs), fuzzy DM, etc. as candidates for DM (see [23] for review on bosonic DM).
@@
Boson stars (BSs), self-gravitating stellar configurations of bosonic particles (see [55, 30] for an updated review on BSs), are a consequence of the existence of such scalar fields in the Universe. These can form from primordial scalar fluctuations, gravothermal collapse in an early matter-dominated era [41], or via a process called gravitational cooling [45]. They can also be formed via trapping of DM in objects like stars and planets [17, 21, 20, 22]. BS is one of the possible exotic compact objects (ECOs) that can mimic other compact objects like black holes (BHs) and neutron stars (NSs) and even source PBHs formation [25]. BSs, if they exist, would also contribute to the total DM budget (see [12] for a review on ECOs) of the universe and their existence would be a key to the DM problem, revolutionizing our understanding of the universe.
@@
Conventional electromagnetic telescopes can probe BSs only if they are accreting matter [35, 44, 43], keeping non-accreting ones out of reach. However, the successful detection of gravitational waves [1] has opened up a new window to probe the dark sectors of the universe. Further, the planning of the next-generation detectors covering a wide frequency range makes the study of all possible gravitational wave (GW) sources of BSs very crucial and timely. There have been considerable efforts to find observational signatures that will be able to distinguish BSs from BHs. Some studies suggested the distinguishability of BSs and BHs as well as NSs based on tidal deformability [11, 46, 24], spin-induced multipole moments [28, 29, 36, 52], II-Love-$Q$ relations [33] and mergers [14].
@@
The study of quasinormal modes (QNMs) is crucial as it carries information about the nature of the underlying compact object [9, 10]. There have been a handful of studies on BS QNMs [26, 56, 6, 32], however, these considered weakly interacting systems. Our focus in this work are massive BSs in the strong-interaction limit (refer to Sec. 2 for details). Flores et al. [54] is the first work to explore $f$-modes, which couple the strongest to GWs, for massive BSs in the strong-interaction limit. The study was restricted to NS mass BSs, and the parameters were restricted by the bounds on self-interaction cross section $0.1<\sigma/m<10$ [cm2/g]. Another work [13], that appeared during the completion of this work, again was restricted to select parameters in the same regime and explored $f$-mode universal relations.
@@
In this letter, we show for the first time that the $f$-mode eigenfrequency follows a scaling relation in the strong-interaction limit, proposing a new signature based on $f$-mode oscillations to delineate massive BSs from BHs and NSs. This scaling is significant, as it allows commenting on the $f$-mode characteristics of BSs with arbitrary self-interacting scalar DM parameters. Using this, we make a detailed comparison of BS $l=2$ QNMs with those of BHs and NSs to comment on the distinguishability. Throughout this work, we use the convention $\hbar=c=1$ and $G=1/M_{Pl}^{2}$, where $M_{Pl}=1.22\times 10^{22}\text{ MeV}=2.18\times 10^{-8}\text{ kg}$ is the Planck mass.
2 Model
@@
For pure Bosonic dark stars, we describe DM by a scalar field ($\phi$) self-interacting via quartic self-interactions as described in [15, 54]. The Lagrangian is given by
@@
where $\lambda$ is the self-interaction parameter and $m$ is the DM particle mass. The self-interaction is negligible only if $\lambda|\phi|^{4}/m^{2}|\phi|^{2}\ll 1$ [15], where $|\phi|\sim M_{Pl}$. A dimensionless parameter is defined as $\Lambda\equiv\lambda M_{Pl}^{2}/4\pi m^{2}$.
For $\Lambda\gg 1$, Colpi et al. [15] showed that massive stars can be formed with mass resembling that of a degenerate fermionic star. In this case, $M_{max}=0.22\sqrt{\Lambda}M_{Pl}^{2}/m=0.062\sqrt{\lambda}M_{pl}^{3}/m^{2}$ [15]. Moreover, $\Lambda\gg 1$, can be achieved for $\lambda\ll 1$. In this limit, even if $\lambda\ll 1$, the interaction is important, and the macroscopic properties are dictated by the strength ($\sqrt{\lambda}/m^{2}$).
@@ 3 Scaling
Colpi et al. [15] showed that the energy density and pressure scale as $\rho^{\prime}=\rho x^{2},~p^{\prime}=px^{2}$ to obtain the scaled version of the equation of state
@@
The dimensionless mass ($M^{\prime}$) and radius ($R^{\prime}$) can then be defined for BSs as $M^{\prime}=M/(xM_{Pl}^{3}),~R^{\prime}=R/(xM_{Pl})$ and can be used to obtain scaled TOV equations.
@@
$f$ $=f^{\prime}/(xM_{Pl}),~\tau=\tau^{\prime}(xM_{Pl})~.$ (7)
We explicitly derive this scaling using the $f$-mode perturbation equations for the case of Cowling approximation, where the background metric is held constant (see A) as well as in full general relativity (see B). This is one of the main results of this work. We now discuss the implications of this scaling.
4 $f$-modes
@@
Figure 1: The scaled $f$-mode frequency ($f^{\prime}=f(xM_{Pl})$) for massive BSs in the strong interaction limit ($\Lambda\gg 1$) as a function of scaled mass $M^{\prime}=M/(xM_{Pl}^{3})$ and compactness $C$. These are unique curves and the scaling relations can be used to obtain the corresponding fMf-M and fCf-C for any microscopic parameters ($\lambda$, $m$). For a given set of parameters, the $f$-mode frequency corresponding to the maximum compactness configuration is the highest and is given by $f_{max}=f^{\prime}_{max}/(xM_{Pl})=0.21/(xM_{Pl})$.
@@
Figure 2: The scaled $f$-mode damping time ($\tau^{\prime}=\tau/(xM_{Pl})$) for massive BSs in the strong interaction limit ($\Lambda\gg 1$) as a function of scaled mass $M^{\prime}=M/(xM_{Pl}^{3})$ and compactness $C$. These are unique curves and the scaling relations can be used to obtain the corresponding $\tau-M$ and $\tau-C$ for any microscopic parameters ($\lambda$, $m$). For a given set of parameters, the $f$-mode damping time corresponding to the maximum compactness configuration is the lowest and is given by $\tau_{min}=\tau^{\prime}_{min}(xM_{Pl})=1900(xM_{Pl})$.
@@
This single solution is sufficient to reproduce all the results and for the entire parameter space in the strong coupling regime ($\Lambda\gg 1$) using the relations $\tau^{\prime}=\tau/(xM_{Pl})$, $M^{\prime}=M/(xM_{Pl}^{3})$ and $C^{\prime}=C$. For a given set of parameters, the minimum possible BS $f$-mode damping time corresponding to the maximum compactness configuration is given by $\tau_{min}=1900\sqrt{\lambda}M_{Pl}/m^{2}$ (see Eq. 7).
@@
5 Comparison with neutron stars and black holes
It would be interesting to compare the BS frequencies with those of the QNMs of other astrophysical compact objects. Here, we compare with the fundamental frequencies of NSs and BHs, which are dominant modes in the context of GW.
@@ 5 Comparison with neutron stars and black holes
$\boxed{f_{BS}=f^{\prime}(C)M^{\prime}(C)\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)=2.03\times 10^{2}f^{\prime}(C)M^{\prime}(C)\left(\frac{M_{\odot}}{M_{BS}}\right)\text{kHz}}~.$ (11)
This is a very useful relation as the frequency can be computed using just two physical parameters, i.e., $M$ and $C$. We can use the unique solutions obtained in Sec. 4 to evaluate $f^{\prime}(C)$ and $M^{\prime}(C)$ (see Fig. 1). The blue regions in Fig. 3 are obtained by varying the compactness for a given mass. The case of maximum compactness (dark blue curve in Fig. 3) is obtained by using $C=C_{max}$, arriving at the equation
$\boxed{f_{BS,max}=0.013\left(\frac{M_{Pl}^{2}}{M_{BS}}\right)=2.6\left(\frac{M_{\odot}}{M_{BS}}\right)\text{kHz}}~.$ (12)
@@
Figure 3: $f$-mode frequency as a function of mass for (a) NS, (b) stellar-mass BH, and (c) SMBH mass range for massive BSs. The blue region is the part where BSs admit a solution. The blue curve corresponds to the case with maximum compactness. The black line is the corresponding BH QNM frequency [27, 7]. The BS frequencies satisfy $f_{BS}/f_{BH}<0.22$ (see text for more details). The silver curves are indicative of the region spanned by NSs [38].
The $f$-mode oscillations of NSs have been studied in great detail. Given the underlying uncertainty in NS EoS, we obtain a band in the $f$-$M$ space. NSs lie in the mass range of 1-2.5 $M_{\odot}$ with compactness in the range 0.1-0.3. We show this mass range in Fig. 3 (a). The blue contour shows the region covered by BSs. The upper boundary corresponds to $C=C_{max}=0.16$. Silver curves show the fMf-M region spanned by NSs due to the uncertainty in the microscopic equation of state (taken from [38]). We see a partial overlap with the NS band, but BS frequencies in the higher mass range are lower than those of NSs. Thus, GW observation of $f$-mode oscillations of a compact object in the NS mass range but outside the expected NS frequency band would be a signature for BS.
Astrophysical BHs are not expected in the mass range of NSs. However, PBHs, if they exist, may have mass in the solar and sub-solar mass range. On the other hand, stellar mass BHs fall in the range 5-100 $M_{\odot}$ and supermassive BHs (SMBH) in galactic centers have masses $10^{6}$-$10^{9}$ $M_{\odot}$. We show these two mass ranges in Figs. 3 (b) and 3 (c), respectively. BHs with masses larger than 100 $M_{\odot}$ are possible via mergers of lighter stellar mass BHs [2]. We show the $n=0$, $l=2$ fundamental $f$-mode QNM frequency [27] by the black curves in Fig. 3. This is given by $\omega M=0.3737-0.0890i$ [27, 7] in geometrical units ($G=c=1$). Reverting to natural units means
$f_{BH}$ $=0.059\left(\frac{M_{Pl}^{2}}{M_{BH}}\right)=1.2\left(\frac{10M_{\odot}}{M_{BH}}\right)\text{kHz}~.$ (13)
The BH frequency only depends on the mass of a non-rotating, chargeless BH (no hair). The BH frequency is also consistently higher than that of BS by a certain factor. This is expected from the high compactness of BH ($C=0.5$), which is higher by a factor of $3.125$ compared to the most compact BS. This factor is readily calculable using Eqs. 12 and 13. We define a ratio ($\mathcal{R_{BS/BH}}$) as the ratio of the fundamental QNM frequency of the most compact BS to that of BH of the same mass. This ratio is a constant, independent of mass, and is given by
@@
The BH frequency is higher than the most compact BS configuration of the same mass throughout by a factor of $0.22^{-1}=4.5$. We thus conclude that BS $f$-modes are always lower than BH frequencies of the same mass by a factor of at least 4.5, i.e., $f_{BS}/f_{BH}<0.22$. If we observe $f$-modes from dark compact objects with a frequency consistently less than $\mathcal{R_{BS/BH}}$ times the expected BH frequency, it would be a smoking gun for massive BSs, which is the main takeaway of this analysis. Note that a similar analysis for other potentials ($V(\phi)$), as well as for another type of BSs, e.g., Oscillatons and Proca Stars, is necessary to distinguish between different models for stars made of Bosonic DM. However, this still acts as a clear signature for distinguishing BHs from BSs. In general, for a BS of arbitrary compactness, the ratio is given by
$\frac{f_{BS}}{f_{BH}}=16.8f^{\prime}(C)M^{\prime}(C)~.$ (15)
We make another comparison with BHs. A relation similar to Eq. 11 exists for the innermost stable circular orbit (ISCO) frequency ($f_{ISCO}$[19]
@@
Here, $M$ is the component mass in an equal mass binary. The ISCO frequency marks the end of the inspiral phase during the binary merger. For $C=0.5$, we get the ISCO frequency for BH. We compare the $f$-mode frequency with the ISCO, as it marks the importance of dynamical tides during pre-merger in the case of a BS binary system. For BHs, the ratio is
$\frac{f_{BH}}{f_{BH}^{ISCO}}=0.3737\times 6^{3/2}=5.5~.$ (17)
Note that this number is not as important for BHs in the context of dynamical tides as it will be for BSs, since there are no BH tidal deformations. For BSs, we get
@@
In the case of BSs, the ratio depends on compactness. The maximum value obtained for the most compact case is 6.6, higher than that for BH. The reason is that although the QNM frequency is higher for BH, the ISCO frequency is much lower in the case of BSs, which are much less compact.
6 Summary